OpenStudy (anonymous):

simplify using radical form: 25 1/2

OpenStudy (anonymous):

if: \[25^{1/2}\] Then: \[=\pm5\]

OpenStudy (anonymous):

like heck

OpenStudy (anonymous):

common cause of confusion. the solution to \[x^2=25\] is \[x=\pm5\] but there is only one square root of 25 and it is 5 and no other

OpenStudy (anonymous):

\[25^{\frac{1}{2}}=\sqrt{25}=5\]

OpenStudy (asadkarim7):

(25)^1\2 (5^2)^1\2 simplify powers 2 * !/2= 1 the remaining is +- 5

OpenStudy (anonymous):

@ satellite73, what you are saying is that the \[\sqrt{25} \neq -5\] and that is false

OpenStudy (anonymous):

lordamercy. once more with feeling. \[\sqrt{x}\] means the POSITIVE number whose square is x. c'est tu

OpenStudy (anonymous):

The PRIME square root of 25 = 5

OpenStudy (anonymous):

in math we only have one answer to the square root of 25. there are two numbers whose square is 25 i agree. they are 5 or -5. so if you are asked to solve \[x^2=25\] you must reply x = 5 or x = -5. that is IF \[x^2=25\] THEN x = 5 or x = -5

OpenStudy (anonymous):

however.... i repeat. the square root of 25 which is what \[25^{\frac{1}{2}}\] means, that is \[25^{\frac{1}{2}}=\sqrt{25}\] denotes not SOME number whose square is 25, but the POSITIVE number whose square is 25 and that number is 5 and no other. the square root function is a perfectly well defined function. it does not give two values. its domain is all numbers greater than or equal to zero and its range is the same. you can bank on it

OpenStudy (anonymous):

hehe... well... just to be that guy.. it says simplify in radical form so the answer should just be \[\sqrt{25}\] ;)